OpenStudy (anonymous):

Photon Lighting Company determines that the supply and demand functions for its most popular lamp are as follows: S(p) = 400 - 4p + 0.00002p4 and D(p) = 2800 - 0.0012p3, where p is the price. Determine the price for which the supply equals the demand.

OpenStudy (anonymous):

Do you know how to star this problem?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

If you were to plot these two, find where they intersect.

OpenStudy (anonymous):

They would intersect at x=96.24 and -118.26. Just chose the positive value.

OpenStudy (anonymous):

That would be were supply meets demand.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

How did they get this answer? :O @zepdrix

zepdrix (zepdrix):

\[\Large\bf\sf S(p) \quad=\quad 400 - 4p + 0.00002p^4\]\[\Large\bf\sf D(p) \quad=\quad 2800 - 0.0012p^3\]And they want to know when:\[\Large\bf\sf S(p)\quad=\quad D(p)\]

zepdrix (zepdrix):

\[\Large\bf\sf 400-4p+0.00002p^4\quad=\quad 2800-0.0012p^3\]Then we solve for p, right? :o

OpenStudy (anonymous):

:O! true true, and then just combine like terms and all? @zepdrix

zepdrix (zepdrix):

Hmm so we don't have any like terms, except for the constants. So moving some stuff around gives us:\[\Large\bf\sf 0.00002p^4+0.0012p^3-4p-2400\quad=\quad0\]

zepdrix (zepdrix):

Hmmm

zepdrix (zepdrix):

Why'd you ask about this question? Did you have the same one come up in your class or something? :D

OpenStudy (anonymous):

I did this practice test (online classes) , and I'm reviewing, and I didn't get this one. So I'm all like :ssssssss

OpenStudy (anonymous):

xD ,So yeah, Like how would I solve this? Like can you help me step by step (with the math)

zepdrix (zepdrix):

Hmm lemme see if I can figure this one out. So we'll multiply by some really big number like 100,000 giving us:\[\Large\bf\sf 2p^4+120p^3-400,000p-240,000,000\quad=\quad0\]

zepdrix (zepdrix):

That get's rid of the decimals. So from here we can apply the Rational Root Theorem.

zepdrix (zepdrix):

Hmm the numbers don't seem to work out nice and neat though :( So I don't think that will work...

zepdrix (zepdrix):

Hmm I dunno how you would solve this one. There isn't a nice fancy "Quadratic Formula" for 4th order equations like this.

OpenStudy (anonymous):

hahahha, Yeah I tried too, but thanks anyways bro!

zepdrix (zepdrix):

:3